OpenStudy (anonymous):

A = { x | x^2 - 4x + 4 = 1 }, B = { 1, 3 }. Show that A = B.

OpenStudy (anonymous):

I have: Let x belong to B. Then x = 1 or 3. 1^2 - 4(1) + 4 = 1. 3^2 - 4(3) = 1. Therefore B is a subset of A. So what about the x in A and A being a subset of B hence A = B?

jimthompson5910 (jim_thompson5910):

Now let y be a number in set A If y is an element of A, then y^2 - 4y + 4 = 1 Solve that for y to get _____ and _______

OpenStudy (zzr0ck3r):

Did you get the last problem?

OpenStudy (zzr0ck3r):

As much as I love spending time on something and getting no reply. At least maybe tell me if it made sense after I make it look all pretty:) @calyne